# Physics Informed Generative Machine Learning for Accelerated Quantum-centric Supercomputing

### v1

@@ I Introduction
Quantum computing holds immense promise for efficiently solving complex and classically intractable problems. A key application area is quantum chemistry, which seeks to determine the ground and excited state energies of many-body fermionic systems by solving the time-independent Schrödinger equation within the Born-Oppenheimer approximation. Designing efficient and accurate quantum algorithms is key toward solving some of the open problems in chemistry such as computation of ligand binding affinity of a drug molecule at a given protein target site[1, 2, 3], accurate electronic structure calculations of the Fe-S clusters for a better understanding of biological nitrogen fixation under ambient conditions[4], or designing high temperature superconducting materials, which involve strong many-body correlation effects driven by Cu 3d electron bands[5]. with the current degree of hardware noise, quantum algorithms like the quantum phase estimation (QPE)[6, 7] is not viable due to its requirement of deep circuits and long coherence time. Consequently, near-term research focuses on hybrid quantum-classical approaches like the variational quantum eigensolver (VQE)[8, 9, 10], quantum approximate optimization algorithm (QAOA)[11] and projective quantum eigensolver (PQE)[12] where quantum processors are dedicated to the evaluation of expectation values and classical processors perform parameter optimization. However, optimization of such hybrid algorithms for practically relevant systems is not guaranteed due to Barren Plateaus[13, 14, 15] and far-from-optimal local traps[16] in the optimization landscape, prolonged optimization runtimes and steep measurement overhead. Reducing the resource requirements for achieving practical quantum advantage of such hybrid algorithms is still an active area of research[17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30]. Despite such intensive research efforts, scaling up electronic structure calculations solely with such variational optimization algorithms remains a nontrivial challenge. In this scenario, alternative algorithmic developments are required to scale up quantum chemical calculations using quantum computers for practical quantum advantage.
@@
Toward this, recently a distinct class of hybrid architecture termed Quantum-Centric Supercomputing (QCSC) has garnered attention due to its enhanced noise resilience and capability of sampling classically hard probability distributions, presenting a potential path forward for larger-scale simulations. QCSC is an emerging paradigm characterized by the seamless integration of quantum computing resources with traditional High-Performance Computing (HPC) infrastructures. This unified architecture is designed to address highly complex, real-world computational problems by distributing the workload to optimally leverage the distinct advantages of both quantum and classical processors. Some of the recently developed QCSC algorithms are inspired by classical selected configuration interaction (SCI)[31, 32, 33, 34, 35] such as quantum SCI (QSCI)[36], sample-based quantum diagonalization (SQD)[37, 38]. In such QCSC methods, generally a state is prepared on a quantum computer that approximately represents the target state and repeated measurements are performed to obtain samples that produce a set of computational basis states. For a molecular system these sampled computational basis states represent important electronic configurations which have dominant contribution to the target electronic state. HPC is required post-sampling for projection of the Hamiltonian in this electronic configuration subspace and diagonalization to get the energy of the approximate state. However, due to the presence of hardware noise, the sampled distributions often contain a substantial number of states that violate fundamental physical symmetries of the molecular system, such as particle number or spin conservation, leading to erroneous results upon diagonalization. To filter these noise-induced configurations and faithfully recover the target state, SQD employs a configuration recovery (CR) module. CR expands the subspace to generate more dominant electronic configurations in a self-consistent manner based on the average occupancy of the corresponding spin-orbitals with subsequent diagonalization. Since sampling dominant configurations from Hilbert space is a computationally challenging task, SQD has the potential to provide advantage over classical SCI methods, provided it produces accurate energy estimations using comparatively reduced dimensional configuration subspace for diagonalization. While SQD with its configuration recovery can produce accurate results for chemically complex systems under strong electronic correlations[37], the accuracy and efficacy depend on the prepared state and the quality of the hardware samples which should have sufficient overlap with the target state for any practical advantage against classical methods. For this hardware-efficient ansatz can be used with parameters optimized in a sub-optimal manner via VQE. However, such an approach adds a heuristic component to the workflow which does not guarantee a substantial overlap with the target state to show any deterministic practical advantage. To this end, recently active research is going on to enhance the performance of SQD to push its limits toward better sampling of the configurations[39, 40]. Given that the state fidelity and overall performance of the SQD pipeline is directly dependent on the quality of this recovery step, the development of a more robust and efficient configuration protocol with better convergence guarantees is critical for the optimal utilization of HPC resources and the scalability of QCSC algorithms as a whole.
@@
In this work, we propose a QCSC workflow driven by a physics-informed generative machine learning (ML) embedded configuration recovery strategy termed PIGen-SQD for accurate fermionic state reconstruction toward improving the efficiency and scalability of SQD algorithm. The development has two primary objectives. First, it provides a physics-guided, deterministic support to the target state via perturbative measures that complement the heuristic state-preparation step, thereby ensuring a substantial initial overlap for the configuration-recovery module. Second, it leverages the ability of generative ML models to learn complex nonlinear distributions in order to reduce the effective diagonalization dimension through classical sampling from the dominant sector of the Hilbert space. To facilitate this generative ML-based configuration recovery, restricted Boltzmann machines (RBM) are used in this work which shows immense promise to solve many-fermionic problems[41, 42, 43, 44, 45, 46]. Our numerical demonstrations on strongly correlated molecular systems indicate that PIGen-SQD reduces the diagonalization subspace dimension by up to 70 relative to standard SQD, while simultaneously yielding energies that are an order of magnitude more accurate. This highlights its ability to reconstruct fermionic states efficiently by exploring only the chemically relevant sector of the Hilbert space. Such a strategic orchestration of classical many-body theory, quantum sampling, and the generative capabilities of machine learning results in an QCSC framework that is accurate, efficient, and systematically improvable.
@@ II Brief Overview of Quantum Centric Supercomputing
where are the composite hole-particle indices. The dimension of space for closed-shell systems is given by with () being the number of electrons in the up (down) spin sectors. Generally the FCI coefficients for some low-lying eigen states are obtained via projection and diagonalization of the Hamiltonian operator in the symmetry space using efficient diagonalization methods like Davidson’s algorithm[47]. However, since the dimension of the symmetry space grows combinatorially with the system size, FCI method cannot be applied to practically relevant systems. Usually space is sparse since a relatively huge number of configurations have zero coefficients and are redundant in the CI expansion Eq. (3). The subspace which contains determinants with only non-zero coefficients is to be referred to as the core space with number of determinants which is in practice much smaller than . Finding out the exact core space is an extremely non-trivial task and the holy grail of any CI method as the prior information of the core space drastically reduces the dimensionality of the diagonalization task for better scalability. Toward this several selected CI methods have been developed recently for selecting subspaces that approximate core space while retaining the desired accuracy[31, 33, 34, 35, 48].
Quantum centric supercomputing (QCSC) or quantum selected CI (QSCI) methods are a class of hybrid quantum-classical algorithms designed to outsource this classically nontrivial core space selection task to a quantum processor This is followed by projection and diagonalization of the corresponding Hamiltonian in the selected space on a classical computer for energy estimation. The fermionic electronic-structure problem is mapped to the quantum-computing framework by encoding the M molecular spin-orbitals onto M qubits, such that each electronic configuration corresponds to a computational-basis state where . The single-qubit states or represents unoccupied and occupied spin-orbitals, respectively. Jordan-Wigner transformation is used to map the corresponding second-quantized fermionic operators to qubit operators. In the QCSC framework, initially a trial state on qubits (which represent molecular spin orbitals) is prepared on a quantum computer with a parameterized quantum circuit (PQC) that should have overlap with the true target state. This is ensured by preparing a state with a physically motivated parameterized ansatz structure of the general form where the unitary operator (ensured by the anti-Hermiticity of the operator ) explicitly introduces electronic corrleation by acting upon the reference state. This PQC is executed on a quantum computer to perform repeated measurements in the computational basis that produce a distribution of measurement outcomes
@@
Here, is the frequency of the configuration in the sample distribution and is the number of shots. In an ideal noiseless scenario the sampled distribution contains the dominant electronic configurations from symmetry space according to the noiseless distribution with the reference configuration being the most frequent. One then diagonalizes the Hamiltonian in the sampled subspace which results in the upper bound energy estimation of the target state. However, in presence of hardware noise, several spurious states contaminate the sampled space with particle or spin non-conserving determinants. This warrants the need for a robust post-processing technique in the QCSC framework that discards the spurious states and recovers the symmetry-preserving configurations in the diagonalization space based on the information from hardware samples, along with exploring additional dominant configurations that the quantum samples might have missed due to hardware noise and inadequate number of shots. The recently developed SQD method introduces a self-consistent configuration recovery protocol to restore the dominant configurations. SQD uses the local unitary coupled jastrow (LUCJ)[49] ansatz which is a variant of the unitary coupled jastrow (UCJ) with a “local” approximation to respect the limited qubit connectivity of a quantum hardware while preserving a degree of physical intuition of the corresponding molecular system. LUCJ with a single layer has the form
@@
where, and are one-body operators and is a density-density operators with being the number operator. In practice, the LUCJ parameters can be initialized with jastrow factorization of the coupled cluster with singles and doubles (CCSD) amplitudes and can be optimized further variationally for a better state preparation[50]. With the hardware samples obtained from quantum circuits with LUCJ approximation, SQD configuration recovery module filters out the spin and particle number symmetry preserving state and partially recovers a new set of configurations from noisy samples with a problem-motivated clustering of configurations using their average orbital occupations. This is followed by an SCI loop for Hamiltonian projection and diagonalization performed in batches each containing configurations sampled from . While the SQD workflow produces accurate results, it is sensitive to the quality of the hardware samples. The presence of huge number of spurious symmetry non-preserving configuration is extremely likely and it requires substantially large number of samples to draw and search for CR module to effectively find out dominant configurations which lie within the core space. This ultimately proliferates the diagonalization dimension which could be the main bottleneck for scalability using HPC resources. To address this issue, in this work, we propose a novel QCSC strategy to restrict the search space and diagonalization dimension of the CR module by providing judicious support to the hardware samples by first-principle many-body perturbative estimates with low rank implicit tensor decompositions and subsequent guided subspace expansion with generative ML.
@@
In the presence of hardware noise and a finite number of shots, even chemistry-inspired hardware-efficient ansatze struggle to generate a sufficient number of samples within the core space which ultimately determine the accuracy of chemically relevant observables. Although such noisy samples can still serve as input for configuration recovery, subsequently this step often requires diagonalizing a large subspace in order to isolate the true core configurations. This diagonalization constitutes one of the primary bottlenecks of QCSC/QSCI workflows and significantly limits scalability. To alleviate this bottleneck, we introduce a first-principles strategy that reliably identifies a dominant configuration subspace from the core space. This subspace, when combined with the symmetry-preserving hardware samples, results in a physics-informed distribution concentrated on configurations with substantial overlap with the target state. Such a distribution provides an excellent initialization for any self-consistent configuration-recovery module, enhancing both the efficiency and the accuracy of subsequent fermionic state reconstruction toward better scalability.
@@ III.1 Anchoring the Hardware Samples to Ground State via Many-Body Theories
Starting from the first principle many-body theory, an unbiased ensemble of the ground state configurations may be built up via Møller-Plesset perturbation theory (MPPT) with the perturbing potential where denotes the Hartree-Fock potential. The corresponding Hilbert space is split into a model space spanned by a reference function and a perturbative orthogonal space spanned by excited determinants defined by . In the perturbative expansion, the exact ground state can be written as
@@
where, with quantities of type being the two-electron integrals and are the first order amplitudes and are generated as a part of the HF module. In the first order correction, only those doubly excited determinants contribute for which is non-zero or in practice, higher than a pre-defined threshold . The same numerical threshold would be used to filter out all the subsequent integrals as well.
@@
Strong correlation often warrants the inclusion of high rank excitations for better representation of the wavefunction. In view of this, we further provide a better support to the ground state wavefunction by identifying the configurations that appear in the second and higher order perturbative rungs. A hierarchical configuration screening ensures that some of the most dominant high-rank configurations are selected to supplement the hardware samples. For the selection of any dominant configuration at a given perturbative order, we analyze in terms of the excitation structure out of the HF vacuum. For any excitation structure (that generates an excited configuration by their action of the vacuum), the many-body perturbation theory ensures that there exists an excitation structure of immediate lower rank which has dominant contribution to the wavefunction. This translates to the fact that the selection of any higher rank excited configuration may be performed by symbolic (rather than numeric) manipulation (or equivalently excitation structure identification) with the knowledge of immediate lower rank configuration.
@@
Nonetheless, since our subsequent workflow does not necessitate such a stringent perturbative measure beyond ensuring adequate support to the target state, we employ an efficient strategy here involving judicious index handling. With the pruned set of and operators, ideally one needs to perform all the tensor contractions of the terms given in Eq. (17) (see Appendix A). Among these terms, computation of is the most expensive as a set of contractible index belongs to the virtual orbital indices. However, due to the sparsity of two body integrals and MP2 amplitudes, this term can be efficiently calculated using a symbolic partial tensor contraction technique for minimal number of operations. Since only the outer indices (non-contractible) post tensor contraction are of importance as the triples are ultimately governed by these outer indices, to reduce the number of operations stemming from explicit tensor contractions, only those combination of outer indices, that originate from pruned and sets, are retained which share only one common index from the contractible set of indices (see Appendix B). Being a symbolic operation, this would not require explicit summation over the entire range of a particular contractible index and the associated cost is in practice much less than . Though such an operation may marginally over-estimate the number of triples and may generate some unimportant configurations, it is guaranteed that the manifold generated by explicit tensor contractions is always a subset of the set of configurations obtained in this manner. Here, the final generated set of triples produced by this operation can also be considered as the set of triples generated in the worst case scenario where all of the screened operators share contractible indices. Moreover, since the nature of the subsequent workflow is agnostic to any such redundant generations (vide infra), such index handling provides the flexibility to avoid the prohibitive cost of explicit tensor contraction, saving huge number of operations to be performed which is further discussed quantitatively in Sec.IV.
@@
At each order of expansion we perform a hierarchical pruning as described above. Since only a small ensemble of configurations contribute significantly to the true wavefunction, these perturbative calculations allow one to efficiently select a substantial amount of the important determinants which can be systematically improved by increasing (along with the associated computational cost). Our numerical analysis shows that the importance selection for configurations obtained from perturbative rank up to provides desired support for the subsequent generative learning which requires asymptotically operations at worst for configuration screening. However, decomposing the corresponding terms (see Appendix A) into intermediates along with the symbolic operation discussed so far, the associated cost would be much reduced. Upon adding the perturbatively obtained configurations to the filtered symmetry-preserved hardware samples, we get a much better distribution of dominant determinants which can be used as an initial point for any self-consistent subspace generation method to predict more determinants from the core space for chemical energy estimation. Toward this, in the next section we review the inner workings of a generative ML method that efficiently scans only an important sector of the symmetry-preserved Hilbert space and further generates important configurations for diagonalization in a self-consistent manner.
III.2 Generative Machine Learning Based Configuration Subspace Expansion
Restricted Boltzmann Machine is a generative stochastic unsupervised machine learning model that has the ability to learn complex nonlinear probability distributions through a bipartite graphical structure consisting of a layer of visible and hidden units. The visible layer encodes the state of the corresponding system and the hidden variables introduce additional degree of freedom for enhanced expressivity. With the training data taken as the input vector for visible layer, RBMs optimize a set of parameters to learn and accurately model the underlying distribution where, are bias vectors for visible and hidden layers and is the weight matrix that connects visible and hidden units. From a mathematical point of view RBMs represent a joint probability distribution over the visible and hidden layers governed by an energy function

### v2

@@ I Introduction
Quantum computing holds immense promise for efficiently solving complex and classically intractable problems. A key application area is quantum chemistry, which seeks to determine the ground and excited state energies of many-body fermionic systems by solving the time-independent Schrödinger equation within the Born-Oppenheimer approximation. Designing efficient and accurate quantum algorithms is key toward solving some of the open problems in chemistry such as computation of ligand binding affinity of a drug molecule at a given protein target site[1, 2, 3], accurate electronic structure calculations of the Fe-S clusters for a better understanding of biological nitrogen fixation under ambient conditions[4], or designing high temperature superconducting materials, which involve strong many-body correlation effects driven by Cu 3d electron bands[5]. However, with the current degree of hardware noise, quantum algorithms like the quantum phase estimation (QPE)[6, 7] is not viable due to its requirement of deep circuits and long coherence time. Consequently, near-term research focuses on hybrid quantum-classical approaches like the variational quantum eigensolver (VQE)[8, 9, 10], quantum approximate optimization algorithm (QAOA)[11] and projective quantum eigensolver (PQE)[12] where quantum processors are dedicated to the evaluation of expectation values and classical processors perform parameter optimization. However, optimization of such hybrid algorithms for practically relevant systems is not guaranteed due to Barren Plateaus[13, 14, 15] and far-from-optimal local traps[16] in the optimization landscape, prolonged optimization runtimes and steep measurement overhead. Reducing the resource requirements for achieving practical quantum advantage of such hybrid algorithms is still an active area of research[17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36]. Despite such intensive research efforts, scaling up electronic structure calculations solely with such variational optimization algorithms remains a nontrivial challenge. In this scenario, alternative algorithmic developments are required to scale up quantum chemical calculations using quantum computers for practical quantum advantage.
@@
Toward this, recently a distinct class of hybrid architecture termed Quantum-Centric Supercomputing (QCSC) has garnered attention due to its enhanced noise resilience and capability of sampling classically hard probability distributions, presenting a potential path forward for larger-scale simulations. QCSC is an emerging paradigm characterized by the seamless integration of quantum computing resources with traditional High-Performance Computing (HPC) infrastructures. This unified architecture is designed to address highly complex, real-world computational problems by distributing the workload to optimally leverage the distinct advantages of both quantum and classical processors. Some of the recently developed QCSC algorithms are inspired by classical selected configuration interaction (SCI)[37, 38, 39, 40, 41] such as quantum SCI (QSCI)[42], sample-based quantum diagonalization (SQD)[43, 44]. In such QCSC methods, generally a state is prepared on a quantum computer that approximately represents the target state and repeated measurements are performed to obtain samples that produce a set of computational basis states. For a molecular system these sampled computational basis states represent important electronic configurations which have dominant contribution to the target electronic state. HPC is required post-sampling for projection of the Hamiltonian in this electronic configuration subspace and diagonalization to get the energy of the approximate state. However, due to the presence of hardware noise, the sampled distributions often contain a substantial number of states that violate fundamental physical symmetries of the molecular system, such as particle number or spin conservation, leading to erroneous results upon diagonalization. To filter these noise-induced configurations and faithfully recover the target state, SQD employs a configuration recovery (CR) module. CR expands the subspace to generate more dominant electronic configurations in a self-consistent manner based on the average occupancy of the corresponding spin-orbitals with subsequent diagonalization. Since sampling dominant configurations from Hilbert space is a computationally challenging task, SQD has the potential to provide advantage over classical SCI methods, provided it produces accurate energy estimations using comparatively reduced dimensional configuration subspace for diagonalization. While SQD with its configuration recovery can produce accurate results for chemically complex systems under strong electronic correlations[43], the accuracy and efficacy depend on the prepared state and the quality of the hardware samples which should have sufficient overlap with the target state for any practical advantage against classical methods. For this, chemistry-inspired hardware-efficient ansatz can be used with parameters optimized in a sub-optimal manner via VQE.[45, 46] However, such an approach adds a heuristic component to the workflow which does not guarantee a substantial overlap with the target state to show any deterministic practical advantage. To this end, recently active research is going on to enhance the performance of SQD to push its limits toward better sampling of the configurations[47, 48]. Given that the state fidelity and overall performance of the SQD pipeline is directly dependent on the quality of this recovery step, the development of a more robust and efficient configuration protocol with better convergence guarantees is critical for the optimal utilization of HPC resources and the scalability of QCSC algorithms as a whole.
@@
In this work, we propose a QCSC workflow driven by a physics-informed generative machine learning (ML) embedded configuration recovery strategy termed PIGen-SQD for accurate fermionic state reconstruction toward improving the efficiency and scalability of SQD algorithm. The development has two primary objectives. First, it provides a physics-guided, deterministic support to the target state via perturbative measures that complement the heuristic state-preparation step, thereby ensuring a substantial initial overlap for the configuration-recovery module. Second, it leverages the ability of generative ML models to learn complex nonlinear distributions in order to reduce the effective diagonalization dimension through classical sampling from the dominant sector of the Hilbert space. To facilitate this generative ML-based configuration recovery, restricted Boltzmann machines (RBM) are used in this work which shows immense promise to solve many-fermionic problems[49, 50, 51, 52, 53, 54]. Our numerical demonstrations on strongly correlated molecular systems indicate that PIGen-SQD reduces the diagonalization subspace dimension by up to 70 relative to standard SQD, while simultaneously yielding energies that are orders-of-magnitude more accurate. This highlights its ability to reconstruct fermionic states efficiently by exploring only the chemically relevant sector of the Hilbert space. Such a strategic orchestration of classical many-body theory, quantum sampling, and the generative capabilities of machine learning results in an QCSC framework that is accurate, efficient, and systematically improvable.
@@ II Brief Overview of Quantum Centric Supercomputing
where are the composite hole-particle indices. The dimension of space for closed-shell systems is given by with () being the number of electrons in the up (down) spin sectors. Generally the FCI coefficients for some low-lying eigen states are obtained via projection and diagonalization of the Hamiltonian operator in the symmetry space using efficient diagonalization methods like Davidson’s algorithm[55]. However, since the dimension of the symmetry space grows combinatorially with the system size, FCI method cannot be applied to practically relevant systems. Usually space is sparse since a relatively huge number of configurations have zero coefficients and are redundant in the CI expansion Eq. (3). The subspace which contains determinants with only non-zero coefficients is to be referred to as the core space with number of determinants which is in practice much smaller than . Finding out the exact core space is an extremely non-trivial task and the holy grail of any CI method as the prior information of the core space drastically reduces the dimensionality of the diagonalization task for better scalability. Toward this, several selected CI methods have been developed recently for selecting subspaces that approximate core space while retaining the desired accuracy[37, 39, 40, 41, 56].
Quantum centric supercomputing (QCSC) or quantum selected CI (QSCI) methods are a class of hybrid quantum-classical algorithms designed to outsource this classically nontrivial core space selection task to a quantum processor. This is followed by projection and diagonalization of the corresponding Hamiltonian in the selected space on a classical computer for energy estimation. Toward this, the fermionic electronic-structure problem is mapped to the quantum-computing framework by encoding the M molecular spin-orbitals onto M qubits, such that each electronic configuration corresponds to a computational-basis state where . The single-qubit states or represents unoccupied and occupied spin-orbitals, respectively. Jordan-Wigner transformation is used to map the corresponding second-quantized fermionic operators to qubit operators. In the QCSC framework, initially a trial state on qubits (which represent molecular spin orbitals) is prepared on a quantum computer with a parameterized quantum circuit (PQC) that should have overlap with the true target state. This is ensured by preparing a state with a physically motivated parameterized ansatz structure of the general form where the unitary operator (ensured by the anti-Hermiticity of the operator ) explicitly introduces electronic corrleation by acting upon the reference state. This PQC is executed on a quantum computer to perform repeated measurements in the computational basis that produce a distribution of measurement outcomes
@@
Here, is the frequency of the configuration in the sample distribution and is the total number of shots. In an ideal noiseless scenario the sampled distribution contains the dominant electronic configurations from symmetry space according to the noiseless distribution with the reference configuration being the most frequent. One then diagonalizes the Hamiltonian in the sampled subspace which results in the upper bound energy estimation of the target state. However, in presence of hardware noise, several spurious states contaminate the sampled space with particle or spin non-conserving determinants. This warrants the need for a robust post-processing technique in the QCSC framework that discards the spurious states and recovers the symmetry-preserving configurations in the diagonalization space based on the information from hardware samples, along with exploring additional dominant configurations that the quantum samples might have missed due to hardware noise and inadequate number of shots. The recently developed SQD[43] method introduces a self-consistent configuration recovery protocol to restore the dominant configurations. SQD uses the local unitary coupled jastrow (LUCJ)[45] ansatz which is a variant of the unitary coupled jastrow (UCJ)[21] with a “local” approximation to respect the limited qubit connectivity of a quantum hardware while preserving a degree of physical intuition of the corresponding molecular system. LUCJ with a single layer has the form
@@
where, and are one-body operators and is a density-density operators with being the number operator. In practice, the LUCJ parameters can be initialized with jastrow factorization of the coupled cluster with singles and doubles (CCSD) amplitudes and can be optimized further variationally for a better state preparation[46]. With the hardware samples obtained from quantum circuits with LUCJ approximation, SQD configuration recovery module filters out the spin and particle number symmetry preserving state and partially recovers a new set of configurations from noisy samples with a problem-motivated clustering of configurations using average orbital occupations. This is followed by an SCI loop for Hamiltonian projection and diagonalization performed in batches each containing configurations sampled from . While the SQD workflow produces accurate results, it is sensitive to the quality of the hardware samples. The presence of huge number of spurious symmetry non-preserving configuration is extremely likely and it requires substantially large number of samples to draw and search for CR module to effectively find out dominant configurations which lie within the core space. This ultimately proliferates the diagonalization dimension which could be the main bottleneck for scalability using HPC resources. To address this issue, in this work, we propose a novel QCSC strategy to restrict the search space and diagonalization dimension of the CR module by providing judicious support to the hardware samples by first-principle many-body perturbative estimates with low rank implicit tensor decompositions and subsequent guided subspace expansion with generative ML.
@@
In the presence of hardware noise and a finite number of shots, even chemistry-inspired hardware-efficient ansatze struggle to generate a sufficient number of samples within the core space[43] which ultimately determine the accuracy of chemically relevant observables. Although such noisy samples can still serve as input for configuration recovery, subsequently this step often requires diagonalizing a large subspace in order to isolate the true core space configurations. This diagonalization constitutes one of the primary bottlenecks of QCSC/QSCI workflows and significantly limits scalability. To alleviate this bottleneck, we introduce a first-principles strategy that reliably identifies a dominant configuration subspace from the core space. This subspace, when combined with the symmetry-preserving hardware samples, results in a physics-informed distribution concentrated on configurations with substantial overlap with the target state. Such a distribution provides an excellent initialization for any self-consistent configuration-recovery module, enhancing both the efficiency and accuracy of subsequent fermionic state reconstruction toward better scalability.
@@ III.1 Anchoring the Hardware Samples to Ground State via Many-Body Theories
Starting from the first principle many-body theory, an unbiased ensemble of the ground state configurations may be built up via Møller-Plesset perturbation theory (MPPT)[57, 58] with the perturbing potential where denotes the Hartree-Fock potential. The corresponding Hilbert space is split into a model space spanned by a reference function and a perturbative orthogonal space spanned by excited determinants defined by . In the perturbative expansion, the exact ground state can be written as
@@
where, with quantities of type being the two-electron integrals, are the second order Møller-Plesset (MP2) amplitudes that are generated as a part of the HF module and are the MP2 denominators.[58] In the first order correction to the wavefunction, only those doubly excited determinants contribute for which is non-zero or in practice, higher than a pre-defined threshold . The same numerical threshold would be used to filter out all the subsequent integrals as well.
@@
Strong correlation often warrants the inclusion of high rank excitations for better representation of the wavefunction. In view of this, we further provide a better support to the ground state wavefunction by identifying the configurations that appear in the second and higher order perturbative rungs. A hierarchical configuration screening ensures that some of the most dominant high-rank configurations are selected to supplement the hardware samples. For the selection of any dominant configuration at a given perturbative order, we analyze in terms of the excitation structure out of the HF vacuum. For any excitation structure (that generates an excited configuration by their action of the vacuum), the many-body perturbation theory ensures that there exists an excitation structure of immediate lower rank which has dominant contribution to the wavefunction. This translates to the fact that the selection of any higher rank excited configuration may be performed by symbolic (rather than numeric) manipulation (or equivalently excitation structure identification) with the knowledge of immediate lower rank configurations.
@@
Nonetheless, since our subsequent workflow does not necessitate such a stringent perturbative measure beyond ensuring adequate support to the target state, we employ an efficient strategy here involving judicious index handling. With the pruned set of and operators, ideally one needs to perform all the tensor contractions of the terms given in Eq. (17) (see Appendix A). Among these terms, computation of is the most expensive as a set of contractible index belongs to the virtual orbital indices. However, due to the sparsity of two body integrals and MP2 amplitudes, this term can be efficiently calculated using a symbolic partial tensor contraction technique for minimal number of operations. Since only the outer indices (non-contractible) post tensor contraction are of importance as the triples are ultimately governed by these outer indices, to reduce the number of operations stemming from explicit tensor contractions, only those combination of outer indices, that originate from pruned and sets, are retained which share only one common index from the contractible set of indices (see Appendix B). Being a symbolic operation, this would not require explicit summation over the entire range of a particular contractible index and the associated cost is in practice much less than . Such index handling provides the flexibility to avoid the prohibitive cost of explicit tensor contractions, saving huge number of operations to be performed which is further discussed quantitatively in Sec.IV.
@@
At each order of expansion we perform a hierarchical pruning as described above. Since only a small ensemble of configurations contribute significantly to the true wavefunction, these perturbative calculations allow one to efficiently select a substantial amount of the important determinants which can be systematically improved by increasing (along with the associated computational cost). Our numerical analysis shows that the importance selection for configurations obtained from perturbative rank up to provides desired support for the subsequent generative learning which requires asymptotically operations at worst for configuration screening. However, decomposing the corresponding terms (see Appendix A) into intermediates along with the symbolic operation discussed so far, the associated cost would be much reduced. Upon adding the perturbatively obtained configurations to the filtered symmetry-preserved hardware samples, we get a much better distribution of dominant determinants which can be used as an initial point for any self-consistent subspace generation method to predict more determinants from the core space for energy estimation. Toward this, in the next section we review the inner workings of a generative ML method that efficiently scans only an important sector of the symmetry-preserved Hilbert space and further generates important configurations for diagonalization in a self-consistent manner.
III.2 Generative Machine Learning Based Fermionic State Reconstruction
Restricted Boltzmann Machine (RBM) is a generative stochastic unsupervised machine learning model that has the ability to learn complex nonlinear probability distributions through a bipartite graphical structure consisting of a layer of visible and hidden units. The visible layer encodes the state of the corresponding system and the hidden variables introduce additional degree of freedom for enhanced expressivity. With the training data taken as the input vector for visible layer, RBMs optimize a set of parameters to learn and accurately model the underlying distribution where, are bias vectors for visible and hidden layers and is the weight matrix that connects visible and hidden units. From a mathematical point of view, RBMs represent a joint probability distribution over the visible and hidden layers governed by an energy function